OpenStudy (anonymous):

Difficult calculus question? A meteorite entering the earth's atmosphere has a velocity inversely proportional to S^0.5 when at a distance of S from the center of the earth. Show that its acceleration is inversely proportional to S^2. I am trying to use the chain rule A=dV/dS x dS/dt but am having problems setting this up.

OpenStudy (anonymous):

any body @lgbasallote @campbell_st want to try!

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

Sorry... Can't help. :( @lgbasallote has done a tutorial on differentiation, perhaps he knows better than me :)

OpenStudy (lgbasallote):

wowww you make a tutorial and suddenly you're supposed to be a master :p haha jk...i dont know this for the reason that i dont know calculus in physics..sorry

OpenStudy (campbell_st):

I thought this may have been of acceleration and velocity in terms of displacement \[\frac{d^2x}{dt^2} = \frac{d}{dx} (\frac{1}{2}v^2)\]